OpenStudy (anonymous):

The table below shows the values of f(n) for different values of n: n 1 2 3 4 5 6 f(n) 1 2 4 16 128 4096 Which recursive function best represents the values shown in the table? A) f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 B) f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2 C)f(1) =1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 D) f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

@welshfella can you help with this one?

OpenStudy (welshfella):

f(1) means the value of function when n = 1 f(2) when n=2 etc

OpenStudy (welshfella):

what you need to do is substitute n = 3 into the formulae for f(n) and see if this gives you f(2) - that is a value of 2.

OpenStudy (anonymous):

oh, ok i'll do that really quickly

OpenStudy (anonymous):

wait:/ i don't have a formula

OpenStudy (welshfella):

sorry - i got that wrong - it should give you the value for f(3) which is 4

OpenStudy (anonymous):

yeah

OpenStudy (welshfella):

ok lets try the last one

OpenStudy (anonymous):

alrighty, what happens first

OpenStudy (welshfella):

f(n) = 2f(n -1) f(n - 2); n > 2 f(3) = 2*f(2)*f(1) = 2*2*1 = 4

OpenStudy (anonymous):

what do you do from there?

OpenStudy (welshfella):

thats the one - we hit the right one

OpenStudy (welshfella):

f(3) = 4 - which is the value on the table

OpenStudy (anonymous):

oh ok

OpenStudy (welshfella):

understand?

OpenStudy (anonymous):

yeah, so just to be clear it would be D?

OpenStudy (welshfella):

yes we could double check to see if f(4) = 16 f(4) = 2*f(3)*f(2) = 2 * 4 * 2 = ?

OpenStudy (welshfella):

does that = 16?

OpenStudy (anonymous):

yes:^)

OpenStudy (welshfella):

thats itt the D is the correct choice

OpenStudy (anonymous):

oh i see, okay thank you @welshfella

OpenStudy (welshfella):

yw